OpenStudy (anonymous):

please help! 4x = 3/7

OpenStudy (ghostgate):

Well, you have to get x on one side of the equation, since you have 4x which is multiplication, what does that leave you? Using the calculator, 3/7 = 0.43 roughly or estimated. So if you have: 4x = 0.43 - What do you do to both sides of the equation to get x alone?

OpenStudy (anonymous):

i don't know, divide?

OpenStudy (ghostgate):

Yes, when you have 4x which is multiplication you would do the 'inverse' as I would say to both sides of the equation, this technically gives you: 1x = ? - 1x can also be seen as simply x, now that you have it alone, what is 0.43 divided by 4 equal?

OpenStudy (ghostgate):

Or 3/7 divided by 4 equal?

OpenStudy (ghostgate):

Well, if you don't know there are two ways of doing this, 1 you do it literally with a calculator or a piece of paper, or you do it alternatively like this: 3/7 / 4 - Change the 4 to a fraction: 3/7 / 4/1 - Alright, since it's division you change things around: 3/7 x 1/4 - Now you even things out by finding the common denominator. 3/7(4) x 1/4(7) - So, now multiply the fractions to get: 12/28 x 7/28 - Now simply multiply: 12*7/28*28 = 84/784 - Now you reduce: 84/784 = 3/28 - So 3/28 is your answer. Although you can also just multiply from here to get the same answer: 3/7 x 1/4 = 3/28 - Either way works. I hope this helps in someway! Have a great day! ;) {---Ghostgate---}